OpenStudy (anonymous):

Find the diameter of a cone that has a volume of 83.74 cubic inches and a height of 5 inches. Use 3.14 for pi.

OpenStudy (anonymous):

my answer was 16, is that correct?

OpenStudy (anonymous):

@mathmale

OpenStudy (harsha19111999):

Volume of a cone = 1/3 * pi * r^2 * h

OpenStudy (anonymous):

well the gave us the volume so i put that in

OpenStudy (harsha19111999):

So you can find out

OpenStudy (anonymous):

i did all my work and my answer was 16

OpenStudy (harsha19111999):

Wait I'll check

OpenStudy (harsha19111999):

16 is r^2

OpenStudy (harsha19111999):

r=4

OpenStudy (anonymous):

can you explain how you got that

OpenStudy (harsha19111999):

So then diameter = 2 * r

OpenStudy (harsha19111999):

\[r^2=16\]